OpenStudy (anonymous):

Find the solution set for (x+7)(x+7)=0

OpenStudy (kidrah69):

\(\huge\color{orange}{WELCOME~TO~OPENSTUDY!!~:)}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (kidrah69):

set each equation to 0 x+7=0 do it once :P

OpenStudy (anonymous):

Hmm.. It said it was wrong. Stupid PLATO

OpenStudy (kidrah69):

eh ?

OpenStudy (anonymous):

PLATO Web. It's a website for school

OpenStudy (kidrah69):

thats a new one :o intresting :) im not sure about this :/ i know that is how u do it

OpenStudy (anonymous):

I hate this website. It tells me my answers are wrong

TheSmartOne (thesmartone):

Welcome to OpenStudy! Did it tell you what the correct answer was then?

OpenStudy (er.mohd.amir):

x=-7 ,-7

TheSmartOne (thesmartone):

And what did you put in as your answer?

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @Er.Mohd.AMIR x=-7 ,-7 \(\color{blue}{\text{End of Quote}}\) That is wrong...

OpenStudy (anonymous):

It was -7

TheSmartOne (thesmartone):

@faketiffanypayne \(\sf\Large (x-a)(x-a)=0\) then: \(\sf \Large x-a=0 \) \(\sf\Large x=a\)

TheSmartOne (thesmartone):

@faketiffanypayne If you did what Pooja told you to do, that is the answer you would have arrived. Also keep in mind that OpenStudy is not meant for giving you answers, we are here to teach you how to get to the answer.

OpenStudy (anonymous):

WEll, yeah. But, that's what the website told me it was

TheSmartOne (thesmartone):

yeah, Pooja or Kidrah told you x+7=0 and if you sovled for x, you were done :)